Can someone help me solve this problem by elimination to come up with the solution?
0.05x + 0.25y=44
0.15x + 0.05y=48

Answer by richwmiller(17219)

x=280 y =120
multiply both by 100
5x + 25y=4400
15x + 5y=4800
multiply first by 3
15x+75y=3*44*100
15x+5y=4800
70y=3*4400-4800
70y=8400
y=120
15x+5y=4800
15x+5*120=4800
15x+600=4800
15x=4200
x=280